Here is the problem. Directions are to solve and check.
3m-(m+1)=6m+1
My work:
1st. 3m-1m-1=6m+1
2nd. 2m-1=6m+1
3rd. 2m-2m-1=-2m+1
4th. -1=4m+1
That is as far as I can get. I dont understand what else to do.

Answer by rapaljer(4671)

You have all the m terms on the right side, so get all the number terms to the left side by subtracting 1 from each side.
-1=4m+1
-1-1=4m+1-1
-2=4m

Divide by 4:
-2/4=4m/4
-1/2=m

R^2